Reports 1-1 of 1 Clear search Modify search
MIF (General)
hirotaka.yuzurihara - 12:56 Monday 14 April 2025 (33395) Print this report
lockloss investigation: 2025/04/08~04/14

I performed the lockloss investigation for the recent lockloss from the OBSERVATION state of the LSC_LOCK guardian, between 2025-04-08 14:51:50.812500 UTC and 2025-04-14 00:03:25.312500 UTC. The previous lockloss investigation was posted in klog33288. During this period, there were 33 lockloss.

Lockloss with the interferometer work

The plots of the 11 lockloss before 4/10 are summarized in wiki. 11 lockloss occurred in the WORKING_WITH_OBSERVATION_WITHOUT_LINES state. (I put the several plots and comments on the wiki)
The plots of the 21 lockloss after 4/10 are summarized in wiki. 6 lockloss occurred in the WORKING_WITH_OBSERVATION_WITHOUT_LINES state. I will summarize the details of rest 15 lockloss below.

1Hz seismic motion made the IMC lockloss

For 9 lockloss, we saw the excess of the seismic motion with 1~10 Hz, which made the oscillation of the IMC length control and made the IMC lockloss.
The time and date are summarized in wiki.

Slowly oscillation (over 5 sec) of OMC DC PD

The aplitude of OMC DCPD increased slowly (over 5 seconds) and saturated, as shown Fig1 and Fig2.
I heard that Yokozawa-san worked at that time and he changed the sign of filter band gain opposite accidentaly. As a result, this oscillation happened (Fig10). We can ignore this lockloss.

  • 1 lockloss: 2025-04-12 22:13:08.687500 UTC

Note that please change the LSC_LOCK guardian state to the WORKING state when you touch the interferometer. Thank you for your cooperation in advance!

Oscillation of ASC (MICH, DHARD, CHARD) at -1 min

At -1 min of the lockloss, the feedback signals for ASC of MICH, DHARD, CHARD looks increasing, as shown Fig3 and Fig4. The oscillation frequency of DHARD was 1.7 Hz.
The oscillation frequency of MICH was ~0.05 Hz.
OMMT2 TRANS increased at -3 s before the lockloss (Fig9).

  • 1 lockloss: 2025-04-13 02:47:00.437500 UTC

Oscillation of ASC with 1.7 Hz

The IR trans and ASC signals of DHARD yaw, CHARD yaw, and CSOFT yaw are oscillating with 1.7 Hz. (Fig5)

  • 1 lockloss: 2025-04-13 04:11:39.812500 UTC

Sudden drift of ASC

At -5 s before the lockloss, the ASC for arm cavities suddenly changed (Fig6). The coincident change can be seen in ASC for IMMT2 in INP2 (Fig7).
Coincidently, AS17 and OMMT2 TRANS increased (Fig8). After that, the IR trans and POP 90 decreased and the lockloss occurred.

  • 1 lockloss: 2025-04-13 04:40:48.187500 UTC

Unknown

For 2 lockloss, there were no clues of the lockloss..

  • 2025-04-11 10:13:18.937500 UTC 
  • 2025-04-13 13:37:30.187500 UTC
Images attached to this report
Search Help
×

Warning

×